OpenStudy (anonymous):

Find the slope- y=(-1/2)x + 1

jimthompson5910 (jim_thompson5910):

y=(-1/2)x + 1 is in the form y = mx+b do you see what the value of m is?

jimthompson5910 (jim_thompson5910):

y = mx+ b m = slope b = y-intercept

OpenStudy (anonymous):

-1/2 is your slope.

OpenStudy (anonymous):

The slope is just 1/2?

jimthompson5910 (jim_thompson5910):

-1/2

jimthompson5910 (jim_thompson5910):

m = -1/2 b = 1

OpenStudy (anonymous):

so if the equation is x - 2y = -8, does that make 2 the slope?

jimthompson5910 (jim_thompson5910):

no you would need to get it into y = mx+b form first

OpenStudy (anonymous):

is it 1/2?

jimthompson5910 (jim_thompson5910):

you are correct, solving for y gives you y = 1/2x + 4

jimthompson5910 (jim_thompson5910):

m = 1/2 b = 4